OpenStudy (anonymous):

if f(x) = 6x + 3 and g(x) = 5x – 4 what is f(g(x))

OpenStudy (anonymous):

f(g(x)) = 6(5x - 4) + 3 f(g(x)) = 30x -24 + 3 f(g(x)) = 30x -21

OpenStudy (anonymous):

Plug in g(x) for the x values in f(x) 6x + 3 6(5x – 4) + 3 30x - 24 + 3 30x - 21

OpenStudy (anonymous):

What f(g(x)) is basically saying, take the y value you get from g(x) and plug that into f(x). Since we are talking generally, the "y value" of g(x) would be 5x - 4 nd plug that whole thing into 6x+3. (plug it into x)